### v1

# ViViDex: Learning Vision-based Dexterous Manipulation from Human Videos

Abstract
In this work, we aim to learn a unified vision-based policy for a multi-fingered robot hand to manipulate different objects in diverse poses. Though prior work has demonstrated that human videos can benefit policy learning, performance improvement has been limited by physically implausible trajectories extracted from videos. Moreover, reliance on privileged object information such as ground-truth object states further limits the applicability in realistic scenarios. To address these limitations, we propose a new framework ViViDex to improve vision-based policy learning from human videos. It first uses reinforcement learning with trajectory guided rewards to train state-based policies for each video, obtaining both visually natural and physically plausible trajectories from the video. We then rollout successful episodes from state-based policies and train a unified visual policy without using any privileged information. A coordinate transformation method is proposed to significantly boost the performance. We evaluate our method on three dexterous manipulation tasks and demonstrate a large improvement over state-of-the-art algorithms. The project page is available at https://zerchen.github.io/projects/vividex.html.
Figure 1: The overall framework of our method for learning dexterous manipulation skills from human videos. It consists of three modules: the extraction of the reference trajectory from human videos, the trajectory-guided state-based policy learning using RL, and the vision-based policy learning using behavior cloning.
I Introduction
Humans possess a remarkable ability to manipulate diverse objects using their hands based on visual perception. Nevertheless, compared with parallel grippers [1, 2, 3, 4, 5, 6], achieving comparable levels of dexterous manipulation with multi-fingered robot hands remains an ongoing challenge [7]. It requires to convert visual signals to high-dimensional control commands.
In recent years, the advancement of deep learning (DL) [8] and reinforcement learning (RL) [9, 10] has led to significant progress in learning-based algorithms for dexterous manipulation such as in-hand rotation [11, 12, 13], solving Rubik’s cube [14] and playing the piano like humans [15, 16]. However, it proves challenging to train RL policies as RL training heavily relies on intricate reward engineering [10, 17] and extensive computational resources [14]. Furthermore, unnatural behaviors with high rewards can emerge from RL training [18]. To address the limitations of RL training, some previous works [19, 20, 21, 22] have turned to imitation learning using robot demonstrations collected by teleoperation. While these approaches enhance training efficiency, they demand substantial human efforts to collect diverse robot demonstration data and are, hence, difficult to scale.
As human videos capturing hands manipulating objects are more abundant and easy to acquire, an increasing research attention has been paid to leveraging human videos for learning dexterous manipulation skills [23, 24, 25, 26, 27]. For example, DexMV [23] proposes to automatically extract robot and object poses from human videos, and then employs the data to accelerate RL via demonstration augmented policy gradient (DAPG) [19]. However, since the extracted robot and object poses are noisy, the method needs hundreds of human videos to learn the manipulation of a single object and still requires reward engineering for different tasks. Additionally, existing works [23, 25, 28] take advantage of privileged information of objects in policy learning such as known object CAD models and ground-truth object poses, which are non-trivial to obtain from raw visual sensors in real-world scenarios.
In this work, we propose a new framework called ViViDex to leverage human Videos for Vision-based Dexterous manipulation policy learning. It consists of three modules as illustrated in Figure 1: reference trajectory extraction, trajectory guided state-based policy learning, and unified vision-based policy learning. We first obtain reference trajectories for hands and objects from human videos following prior work [23]. Though such trajectories are noisy and not directly suitable for robot control, they provide examples of natural hand-object interactions that we use as guidance. Our second module refines and recovers physics properties of the reference trajectory. We use RL for this purpose by training a state-based policy for each reference trajectory. We design a novel reward to keep pose similarity in the reference trajectory while solving the task. We also augment the trajectory during training to generalize to a broader range of object poses beyond the pose in the reference trajectory. Finally, we rollout episodes from optimized state-based policies and distill successful episodes to a unified vision-based policy without any privileged object information. A coordinate transformation method is proposed to improve the vision-based policy, which transforms the input 3D point clouds into hand joints and the target coordinate systems to capture fine-grained interaction features and increase awareness of the target position. We validate the effectiveness of our system on three dexterous manipulation tasks relocation, pouring and placing inside. Our ViViDex approach significantly outperforms the state-of-the-art method DexMV [23] while using significantly less human demonstrations.
To summarize, our contributions are three-fold:
$\bullet$ We introduce a new framework ViViDex to learn vision-based dexterous manipulation policy from human videos.
$\bullet$ We improve the RL rewards for the state-based policy by imitating trajectories extracted from human videos and propose a novel model architecture for the vision-based policy.
$\bullet$ Experiments on three dexterous manipulation tasks demonstrate the effectiveness of the ViViDex approach.
II Related Work
Dexterous manipulation. The multi-fingered robotic hands empower robots to perform delicate operations on objects like humans. Previous works can be generally grouped into two categories: trajectory optimization methods [29, 30, 31, 32, 33] and data-driven learning methods [19, 23, 24, 28, 34, 35, 36, 37]. Trajectory optimization methods usually require well-defined dynamic models for the robot and the manipulated object, which are usually not easy to obtain in practice. Data-driven methods instead train neural network policies for manipulation, allowing to learn generalizable skills. Most of the works [19, 28, 35, 38, 39, 40] use reinforcement learning (RL) algorithms to train policies. Nagabandi et al. [39] and Hansen et al. [40] propose to improve the sample efficiency in RL by jointly learning the dynamic model and the control policy. To accelerate the convergence of RL, Rajeswaran et al. [19] introduce the demo augmented policy gradient algorithm (DAPG) and train RL with expert demonstrations. Instead of manually collecting expert demonstrations, ILAD [28] derives demonstration data via an affordance model [41] and trajectory optimization [42]. Xu et al. [35] and Wan et al. [38] decompose the dexterous grasping into two sub-tasks: the grasp proposal generation and the goal-conditioned policy training and largely improve the generalization of the learned policy. However, RL is data-inefficient and requires well-designed rewards, which often results in unrealistic object manipulation actions. To address these problems, more recent works [23, 25, 26, 27] attempt to learn policies using human demonstrations and combine imitation learning with RL.
Learning robotic manipulation from human videos. Human videos naturally provide rich human hands interactions with diverse objects [43, 44, 45, 46, 47]. Equipping robots with the capacity to acquire manipulation skills by simply watching human videos has been an attractive direction [48, 49, 50, 51, 52]. Some works [34, 53, 54, 55] propose to use human videos to pre-train a visual representation for the policy learning. R3M [34] proposes to pre-train CNNs with egocentric human videos [56] and perform the imitation learning based on the learned visual representations. Xiao et al. [53] and Radosavovic et al. [54] propose to leverage in-the-wild human videos to pre-train the vision transformer and demonstrate the improvement on different manipulation tasks. The recent work H-InDex [55] employs 3D human pose estimation from human videos as a proxy task to pre-train the visual representation for dexterous manipulation tasks. Another line of works [57, 58, 59, 60] focus on learning the reward function from human videos. Chen et al. [57] introduce DVD, a domain-agnostic video discriminator, to learn multi-task reward functions. Alakuijala et al. [58] propose a task-agnostic reward function by using unlabeled human videos. Chane-Sane et al. [59] formulate the reward learning as a retrieval task and propose a data-driven approach to map human demonstrations to unseen robotic manipulation tasks. Xiong et al. [60] propose to translate human demonstration videos into robot demonstration videos and use the keypoint-based representation to define the distance metric for the policy execution. Some other works [23, 24, 25, 26, 61] explore how to explicitly leverage human videos to facilitate learning dexterous manipulation skills. Mandikal et al. [61, 24] propose to utilize object affordances priors and human grasping priors to improve the dexterous grasping policy. Shaw et al. [26] propose to first pre-train the policy with visual, action and physical priors learned from human videos and finetune the policy with teleoperated demonstrations to make it generalize to different robotic tasks. DexMV [23] and DexRepNet [25] extracts demonstrations data from human videos and uses DAPG algorithm [19] to launch the RL training with demonstrations. In this work, not limited to the simple dexterous grasping action [24, 25, 35, 38, 61], we aim to solve more complex tasks by utilizing human videos. To additionally address limitations of previous works [23, 25, 28] that require object states (e.g., object poses) during the policy execution, we propose learning vision-based policies that achieve high manipulation accuracy while only using proprioceptive robot information and depth observations in the form of point clouds.
III Method
TABLE I: The state definition. The left part summarizes robot states. The right part defines object states.
Parameters Descriptions Parameters Descriptions
$\bm{q}_{r}\in\mathbb{R}^{30}$ robot motor positions $\bm{q}_{o}\in\mathbb{R}^{6}$ object canonical positions
$\bm{\dot{q}}_{r}\in\mathbb{R}^{30}$ robot motor velocities $\bm{\dot{q}}_{o}\in\mathbb{R}^{6}$ object canonical velocities
$\bm{x}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot joint positions $\bm{x}_{o}\in\mathbb{R}^{3}$ object center position
$\bm{\theta}_{\text{r}}\in\mathbb{R}^{16\times 4}$ robot joint rotations $\bm{\theta}_{o}\in\mathbb{R}^{4}$ object global rotation
$\bm{\dot{x}}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot linear velocities $\bm{\dot{x}}_{\text{o}}\in\mathbb{R}^{3}$ object linear velocity
$\bm{\omega}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot angular velocities $\bm{\omega}_{\text{o}}\in\mathbb{R}^{3}$ object angular velocity
$\bm{t}_{\text{r}}\in\mathbb{R}^{7\times 3}$ robot target translations $\bm{t}_{\text{o}}\in\mathbb{R}^{3}$ object target translations
This work aims to train vision-based dexterous manipulation policies. However, jointly learning the visual representation and the control policy is challenging. Therefore, we first train the state-based policy and then distill the generated data to a vision-based policy. To train the state-based policy, as the pure RL training is ineffective [23, 25, 28], we extract reference trajectories from human videos and propose trajectory-guided RL to improve the performance. In the following, we first present details about the reference trajectory generation in Section III-A. Then, we introduce our state-based policy learning algorithm in Section III-B. Finally, we describe how we distill multiple state-based policies into the visual policy and its model architecture in Section III-C.
III-A Reference Trajectory Extraction
In this work, we use the DexYCB [43] video dataset, which captures human hand grasping of objects. We first extract hand and object poses from the video. The hand poses and shapes are represented by MANO [62] and result in 3D locations of joints $\bm{\psi}_{h}\in\mathbb{R}^{21\times 3}$. Then, we aim to retarget the human hand pose to the robot hand pose. We use the Adroit anthropomorphic robot [19], which consists of 30 degrees of freedom $\bm{q}_{r}\in\mathbb{R}^{30}$. Following [23, 63, 64], we formulate the hand motion retargeting from a video of length $T$ as an optimization problem and define its objective function as:
$\begin{split}&\min_{\bm{q}_{r}^{t}}\sum_{t=1}^{T}\Big{\|}\bm{\hat{x}}_{rt}^{t}(\bm{q}_{r}^{t})-\bm{{\psi}}_{ht}^{t}\Big{\|}^{2}_{2}\\ &+\Big{\|}\bm{\hat{x}}_{rp}^{t}(\bm{q}_{r}^{t})-\bm{{\psi}}_{hp}^{t}\Big{\|}^{2}_{2}+\alpha\Big{\|}\bm{q}_{r}^{t}-\bm{q}_{r}^{t-1}\Big{\|}^{2}_{2},\end{split}$ (1)
where $\bm{{\psi}}_{ht}^{t}\in\mathbb{R}^{5\times 3}$ and $\bm{{\psi}}_{hp}^{t}\in\mathbb{R}^{5\times 3}$ denote human hand tip and middle phalanx positions, respectively. We solve their counterparts $\bm{\hat{x}}_{rt}^{t}$ and $\bm{\hat{x}}_{rp}^{t}$ in the robot via forward kinematics according to given ${\bm{q}_{r}^{t}}$. In the first two terms, we aim to minimize the difference between the 3D locations of robot joints to those in the human hand. We also use a regularization term to avoid sudden changes in the robot pose, where we empirically set $\alpha$ and ${\bm{q}_{r}^{0}}$ to 4e-3 and the mean pose of its motion limits, respectively. We use the NLopt solver [65] for the optimization and port robot and object motions into the simulator to generate reference trajectories, which are only visually plausible but not physically plausible.
Figure 2: The illustration of different stages for the reference trajectory. The pre-grasp stage (a) describes how the robot approaches the object. The manipulation stage (b) focuses on how the robot hand manipulates the object.
III-B State-based Policy
We employ the generated reference trajectory as the reward function for RL to constrain the robot hand and object motions and train a state-based policy to recover physically plausible trajectories. The network architecture for the state-based policy consists of actor and critic MLPs [66]. It takes both robot and object states described in Table I as inputs and predicts the robot control commands. Different from DexMV [23] that needs different reward functions for different manipulation tasks, our method can solve different tasks with the same reward functions. In the following, we will describe our reward functions and the trajectory augmentation for RL.
Reward functions. To define the reward function, we divide the reference trajectory into two stages shown in Figure 2: the pre-grasp stage and the manipulation stage. The recent work PGDM [27] focuses on the manipulation stage and initializes the robot hand at a given pre-grasp configuration for RL, which is a simplified setting compared with ours.
During the pre-grasp stage, the robot hand approaches the object without making physical contacts. During RL training, we require that the robot approaches the object in a similar way as humans and define the following reward function:
$\begin{split}R_{p}=\sum_{t=1}^{T_{p}}10\cdot{\rm exp}(-10\cdot\Big{\|}\bm{{x}}_{rt}^{t}(\bm{q}_{r}^{t})-\bm{\hat{x}}_{rt}^{t}\Big{\|}^{2}_{2}),\end{split}$ (2)
where $T_{p}$ is the length of pre-grasp steps, and $\bm{\hat{x}}_{rt}^{t}\in\mathbb{R}^{5\times 3}$ denotes the robot finger tip positions at the timestep $t$ in the reference trajectory $\bm{{x}}_{rt}^{t}$ is the current robot tip locations.
When the robot successfully reaches its pre-grasp configuration, the episode starts its manipulation stage. In this stage, the robot manipulates the object and aims to bring it to the desired target configuration. Here, we define the reward function to constrain the robot and object motions jointly:
$\begin{split}&R_{m}=\sum_{t=T_{p}+1}^{T_{r}}\lambda_{1}\cdot{\rm exp}(-\alpha_{1}\cdot\Big{\|}\bm{{x}}_{rt}^{t}(\bm{q}_{r}^{t})-\bm{\hat{x}}_{rt}^{t}\Big{\|}^{2}_{2})\\ &+\lambda_{2}\cdot{\rm exp}(-\alpha_{2}\cdot(\Big{\|}\bm{x}_{o}^{t}-\bm{\hat{x}}_{o}^{t}\Big{\|}^{2}_{2}+\alpha_{3}\cdot\phi(\bm{\theta}_{o}^{t},~{}\bm{\hat{\theta}}_{o}^{t})))\\ &+\sum_{i=1}^{5}\lambda_{3}\cdot\mathds{1}^{\rm contact}_{\rm finger~{}i}+\lambda_{4}\cdot\mathds{1}^{\rm lift}_{\rm object},\end{split}$ (3)
where $T_{r}$ refers to the length of the whole reference trajectory. The first term of (3) constrains the hand motions similar to (2). The second term constrains the object motions, and $\bm{x}_{o}^{t}$ and $\bm{\hat{x}}_{o}^{t}$ represent the current object position and the reference object position at timestep $t$. $\phi(\cdot)$ computes the angular distance between the current object orientation with its reference $\bm{\hat{\theta}}_{o}^{t}$. In the third term, we encourage contacts between each finger of the robot and the object. The forth term assigns bonus points when the object is lifted off the table. The second and forth terms are introduced in [27], and we use the same hyper-parameters (i.e., $\lambda_{2}=10$, $\alpha_{2}=50$, $\alpha_{3}=0.1$ and $\lambda_{4}=2.5$). For our proposed first and third terms, we empirically set $\lambda_{1}$, $\alpha_{1}$ and $\lambda_{3}$ to 4, 10 and 0.5, respectively.
When the robot and object reaches their targets, we want them to remain stable in their target configurations for a while, which is especially important for the pour task to make sure that all particles in the mug have enough time to drop into the target container. To achieve this when $t~{}>~{}T_{r}$, we still use (3) and constrain the robot and object to stay in their final target configurations until the end of the episode.
Figure 3: Illustration of our vision-based policy architecture.
Reference trajectory augmentation. By using our proposed reward functions, though the state-based policy can successfully imitate the hand and object motions from a video, it still cannot generalize to different initial object positions, rotations and target positions. To make our policy applicable to different initial object configurations and target positions, we introduce our reference trajectory augmentation strategies during RL training. To deal with different initial object positions or rotations, we could accordingly transform the whole reference trajectory. When we randomly set the target object position, we augment the original object trajectory through interpolations aimed at reaching the specified target position. For pour and place inside tasks, we first design the desired final object configuration. In the pour task, we ensure that the mug is tilted appropriately, which allows the particles inside to effortlessly drop into the target container. Our goal for place inside is to accurately position the banana within the container. Then, we interpolate the object position and orientation to reach the desired configuration. Similarly, we can also interpolate hand motions and use our proposed reward functions to optimize the policy.
III-C The Vision-based Policy
Though our state-based policy can work well for different manipulation tasks, it requires robot proprioceptive states and object states as inputs. However, reliably estimating object states is often not trivial in practice. To alleviate this issue, we propose a vision-based policy that only takes robot states and 3D scene point clouds as inputs. To generate training data for the visual policy, we rollout the optimized state-based policy and generate diverse physically plausible trajectories. During the rollout process, we compute the 3D point clouds ${{\rm PC_{w}}\in\mathbb{R}^{N\times 3}}$ from three depth cameras in the simulator, where $\rm PC_{w}$ is represented under the world coordinate system (i.e., the center of the table) and $N$ is the number of points.
Figure 4: Motion retargeting results for selected DexYCB [43] videos. The three rows show different initial object poses for each object category.
As shown in Figure 3, inspired by [5], we propose to transform $\rm PC_{w}$ into the desired target coordinate system $\rm PC_{t}$, which make the model more aware of the target position for control predictions. To capture dense interaction features between the robot and the object, we additionally transform $\rm PC_{w}$ into six different robot joint coordinate systems (i.e., one wrist and five fingers). Finally, we combine point clouds representations under different coordinate systems ${\rm PC}\in\mathbb{R}^{N\times 24}$ and feed them into PointNet [67]. Our visual policy predicts control commands based on extracted visual features and robot states. We train our visual policy using behavior cloning from generated physically plausible trajectories.
IV Experiments
We conduct extensive experiments to evaluate our state-based and visual policies on three manipulation tasks, namely relocate, pour and place inside tasks.
IV-A Experimental Setting
Video dataset. The DexYCB [43] dataset contains human demonstration videos of hand-object interactions for 20 YCB objects with 20 videos per object. Following [23], we focus on five objects including mustard bottle, tomato soup can, sugar box, large clamp and mug. Due to the unavailability of DexMV videos [23], we select videos from the DexYCB dataset that shares high visual similarities with DexMV videos. For each object, we choose one video which is most similar to those used in DexMV and use this video for most of our experiments; the top row in Figure 4 presents motion retargeting results for these videos. For the final experiment, we add additional initial configurations of the objects to the test set (different from [23]) and use two additional videos per object with different initial object positions for training, as illustrated in the last two rows of Figure 4. We use hand and object poses provided by the dataset to perform motion retargeting and generate reference trajectories.
Simulation benchmark. We follow previous work [23, 27] and use the MuJoCo physics simulator [68] together with the Adroit robot hand [19]. The robot hand consists of 30 degrees of freedom (DoF). We set the control frequency and simulation frequency as 40HZ and 400HZ, respectively. As discussed above, we use the same testing configuration as in [23] except for the final experiment, where we use a more difficult setting, and evaluate the performance of our method on three manipulation tasks:
TABLE II: Performance of our state-based policy using hand rewards at different stages. It follows the human hand trajectory extracted from videos to manipulate objects (R1), reach the pre-grasp configurations (R2) or combine both (R3).
Hand Reward ${\rm E_{o}}\downarrow$ ${\rm E_{h}}\downarrow$ ${\rm SR_{o}}\uparrow$ ${\rm SR_{h}}\uparrow$ ${\rm SR_{3}\uparrow}$
Pre-grasp Manipulation
R1 $\times$ 0.063 0.15 0.37 0.00 0.00
R2 $\times$ 0.0031 0.053 0.93 0.32 1.00
R3 0.0031 0.024 0.95 0.83 1.00
Relocate task requires the robot to move an object to a target position. The task is successfully solved when the object reaches the target position within a certain threshold.
Pour task requires the robot to grasp a mug filled with particles and pour the particles into a container. The success rate is measured by the percentage of particles in the container.
Place inside task requires the robot to grasp a banana and place it into a mug. The success rate is measured by the percentage of the banana mesh in the mug.
Evaluation metrics. The success rate (${\rm SR}$) is the major metric for the above three tasks. For the relocate task, we adopt the 10cm threshold (${\rm SR_{10}}$) following [23, 28] to define the success rate. To measure the accuracy more precisely, we add a more rigorous threshold of 3cm and compute the success rate ${\rm SR_{3}}$. For the state-based policy, we additionally evaluate how well the learned physically plausible trajectory matches the reference trajectory in the video using four metrics [27]: ${\rm E_{o}}$ computes the average object position error between the current object trajectory and its reference trajectory over time; ${\rm E_{h}}$ computes the average finger tips position error between the current robot hand trajectory and its reference trajectory over time; ${\rm SR_{o}}$ reports the fraction of timesteps where ${\rm E_{o}}$ is below the 1cm threshold; ${\rm SR_{h}}$ computes the fraction of timesteps where ${\rm E_{h}}$ is lower than 5cm.
Implementation details. During RL training, we randomly set the object initial and target positions following [23] and also randomize the initial object orientation. During testing, we run 300 episodes with different initial configurations and report the average success rate. For the state-based policy, we use both robot and object states defined in Table I as inputs and predict the control command. The length of a single episode is 60, 80 and 100 for the relocate task, place inside task and pour task, respectively. We use PPO [66] to optimize the state-based policy. We set the learning rate and batch size to 1e-5 and 256, respectively. The RL training takes around 30 hours on a single V100 GPU and 32 CPUs for each video separately. Our visual policy only requires robot proprioceptive states (the left part of Table I) and point clouds as inputs. To train the visual policy, we rollout 500 successful trajectories from the state-based policy for each video, where trajectories differ in initial object positions, rotations and target positions. During rollouts, we render 3D scene point clouds from three depth cameras surrounding the table in the simulation platform. We train the visual policy with Adam optimizer using a learning rate of 1e-5 and a batch size of 64.It takes 8 hours on 4 A100 GPUs for our visual policy trained from 5 videos.
IV-B Evaluation of the State-based Policy on the Relocate Task
We adopt the relocate task to evaluate the state-based policy. Specifically, we use one DexYCB video for each of the five object categories and train a state-based policy for each video separately. The initial object poses of the videos are illustrated in the first row of Figure 4.
TABLE III: Comparison with state-of-the-art methods on the relocate task for state-based policies. We optimize a state-based policy using a single human video for each object (S6, S7), while previous methods usually need about one hundred human videos for each object (S2-S4). * indicates that we re-train the method from [23] with 20 DexYCB videos for each object.
Methods Training algorithms Num. videos Mustard bottle Tomato can Sugar box Large clamp Mug Avg.
${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
S1 DexMV [23] TRPO [69] 0 0.06 - 0.67 - 0.00 - 0.51 - 0.49 - 0.35 -
S2 DexMV [23] SOIL [70] 97 0.33 - 0.98 - 0.67 - 0.89 - 0.71 - 0.72 -
S3 DexMV [23] GAIL+ [71] 97 0.06 - 0.66 - 0.00 - 0.52 - 0.53 - 0.50 -
S4 DexMV [23] DAPG [19] 97 0.93 - 1.00 - 0.00 - 1.00 - 1.00 - 0.79 -
S5 DexMV [23] DAPG [19] 20 1.00 0.63 1.00 0.56 0.09 0.00 0.08 0.00 0.02 0.00 0.44 0.24
S6 Ours / wo rot. PPO [66] 1 0.75 0.03 0.74 0.65 0.38 0.08 0.39 0.03 0.98 0.03 0.65 0.16
S7 Ours PPO [66] 1 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
TABLE IV: Performance of separate visual policies for each object on the relocate task. The initial configurations in testing are derived from the first row of Figure 4 for each object. Our visual policies (V4-V6) only utilize one video in training. “Comp. PC” denotes complete point clouds sampled from object CAD models without occlusion.
Source Models Comp. PC Object states Visual inputs Num. Points Noisy Points Mustard bottle Tomato can Sugar box Large clamp Mug Avg.
${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
V1 S4 512 $\times$ 1.00 0.67 1.00 0.50 0.14 0.00 1.00 0.77 1.00 0.58 0.83 0.50
V2 S4 $\times$ 512 $\times$ 1.00 0.62 1.00 0.48 0.14 0.00 1.00 0.76 0.99 0.55 0.83 0.48
V3 S4 $\times$ $\times$ 512 $\times$ 0.98 0.55 1.00 0.46 0.14 0.00 1.00 0.76 0.99 0.54 0.82 0.46
V4 S7 $\times$ $\times$ 512 $\times$ 1.00 0.89 1.00 0.86 1.00 0.96 1.00 0.91 1.00 1.00 1.00 0.92
V5 S7 $\times$ $\times$ 2048 $\times$ 1.00 0.97 1.00 0.80 1.00 0.98 1.00 0.93 1.00 1.00 1.00 0.94
V6 S7 $\times$ $\times$ 2048 0.97 0.77 0.97 0.57 1.00 0.88 1.00 0.89 1.00 1.00 0.99 0.82
Figure 5: Qualitative comparison of state-based policies using different rewards. R1 (w/o hand reward in pre-grasp) leads to unstable grasps. R2 (w/o hand reward in manipulation) results in unnatural hand actions. Our proposed policy R3 uses hand rewards at both stages and achieves the best performance.
IV-B1 Ablation Studies
We first perform ablations to validate the importance of our hand rewards at the pre-grasp and the manipulation stages in state-based policy learning. Table II presents the averaged performance over the five videos for policies with different reward functions. The policy in R1 uses the proposed hand rewards (i.e., Equation 3) in the manipulation stage but follows the previous reward function [23, 28, 19] in the pre-grasp stage which is to minimize the distance between the robot hand and the object. It achieves poor performance due to the lack of guidance from human hand trajectory to approach the object. As shown in the first column of Figure 5, the policy is difficult to arrive at a plausible pre-grasp configuration to stably lift the object. The policy in R2 employs the proposed reward function in the pre-grasping stage (Equation 2) but follows PGDM [27] (i.e., the second and forth terms in Equation 3) in the object manipulation stage. The reward function in PGDM constrains the object motions to be close to the reference trajectory but does not add any constraints on the hand motions, which results in unnatural robot hand actions as shown in the second column of Figure 5. Our proposed policy in R3 utilizes hand rewards in both stages, which outperforms R1 and R2 on all the metrics and produces robust and realistic grasps as shown in the last column of Figure 5.
IV-B2 State-of-the-art comparison
Table III presents the quantitative comparison of our state-based policy and state-of-the-art models [23]. S1 to S4 denote four model variants reported in [23]. As their training data is not released, we re-train their best model DAPG [19] (S4) using DexYCB videos and report the results in S5 for more fair comparison. We can see that the performance drops significantly for some objects, i.e., large clamp and mug. This indicates that their method requires sufficient amount of videos to learn the manipulation of objects with complex shapes. Our proposed policy in S7 significantly outperforms the previous methods despite only using one training video per object. Though S4 also performs well for mustard bottle, tomato soup can, large clamp and mug, it easily knocks down thin objects (e.g., sugar box) when the robot approaches the object and achieves poor performance. Our hand reward addresses this problem by mimicking the human pre-grasping trajectory. S6 is a variant of our model without rotation augmentation in training. The inferior performance of S6 to our final model S7 demonstrates the importance of augmentation to help improve the robustness of trained policies.
IV-C Evaluation of the Visual Policy on the Relocate Task
Here, we evaluate our proposed visual policy on the relocate task and present a detailed experimental analysis.
IV-C1 Learning from a single video per object
We first explore training separate visual policies for each object with only one video as in the previous section IV-B. During testing, we use the same object configurations as in the section IV-B, which are derived from object poses in the first row of Figure 4 by randomly changing the initial object position and rotation around z-axis for each testing episodes. The results are presented in Table IV. From V1 to V3, we rollout trajectories from the S4 models in Table III  [^1] We use S4 models available at released codes of DexMV [23]. to train the visual policy. V1 has the same assumption as [28] which assumes known object states and complete object point clouds. V2 removes the dependency on complete point clouds but still requires object states. V3 further relaxes the assumption and only uses robot states together with scene point clouds. We can see that performance decreases when using no privileged object information (i.e., object states and complete object point clouds). From V4 to V6, we train the visual policy using rollouts from our state-based policies S7. Benefiting from high-quality trajectories generated by S7, our visual policy V4 can outperform V1 to V3 by a large margin without privileged information. To investigate how the density of point clouds affect the visual policy, we sample 2048 points in V5 at each timestep. This marginally improves the performance compared to V4. In V6 we evaluate the robustness of our visual policy by adding Gaussian noise with $\sigma$=3cm to 3D point clouds during training and testing. Our visual policy still functions reasonably under noisy point clouds observations.
TABLE V: Performance of separate policies for each object on the relocate task. The initial configurations in testing are derived from all three rows in Figure 4. We compare our visual policies using different number of training videos. “Mb.”, “Tc.”, “Sb.” and “Lc.” denote mustard bottle, tomato can, sugar box and large clamp respectively.
Num. videos Mb. Tc. Sb. Lc. Mug Avg.
${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
S4 97 0.46 0.28 0.59 0.35 0.11 0.00 0.63 0.31 0.65 0.62 0.49 0.31
V5 1 0.33 0.32 0.41 0.25 0.33 0.32 0.43 0.31 0.58 0.57 0.42 0.35
V7 2 0.65 0.42 0.66 0.60 0.66 0.62 0.67 0.64 0.91 0.85 0.71 0.63
V8 3 0.96 0.66 0.95 0.82 0.89 0.63 0.98 0.92 0.98 0.97 0.95 0.80
TABLE VI: Performance of unified visual policies for all objects on the relocate task. The initial configurations in testing are derived from all three rows in Figure 4. We compare point cloud representations with different coordinate systems where “wo.”, “ta.” and “ha.” represent world, target and robot hand. “Mb.”, “Tc.”, “Sb.” and “Lc.” denote mustard bottle, tomato can, sugar box and large clamp respectively.
Coor. sys Mb. Tc. Sb. Lc. Mug Avg.
wo. ta. ha. ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
V9 $\times$ $\times$ 0.77 0.32 0.73 0.44 0.80 0.29 0.87 0.40 0.79 0.58 0.79 0.41
V10 $\times$ 0.91 0.84 0.94 0.92 0.99 0.97 0.96 0.93 0.92 0.92 0.94 0.92
V11 0.94 0.86 0.99 0.98 0.98 0.93 0.99 0.97 0.96 0.95 0.97 0.94
IV-C2 Learning from multiple videos per object
We next introduce a more challenging testing scenario where we evaluate policies under three different initial object configurations for each object as shown in Figure 4. For example, the mustard bottle does not necessarily stand on the table, but can lie in different ways. Table V presents the results of different models. The state-based policy S4 [23] in Table III though employs privilege object information and multiple videos per object in training, still generalizes poorly to different initial object poses. Our visual policy V5 in Table IV also suffers in this challenging testing scenario due to the insufficient training videos. From V7 to V8, we gradually increase the number of training videos with diverse initial object configurations and demonstrate significant increase in performance.
Figure 6: Qualitative illustration of the performance of our visual policies on pour (a) and place inside (b) tasks.
TABLE VII: Comparison with state-of-the-art methods on the pour and the place inside tasks for our policy models. Our approach uses a single human video for each task, while previous methods need more than 91 human videos.
Methods Training algorithms Num. videos Pour success rate Place inside success rate
S1 DexMV [23] TRPO [69] 0 / 0 0.01 0.03
S2 DexMV [23] SOIL [70] 101 / 91 0.04 0.28
S3 DexMV [23] GAIL+ [71] 101 / 91 0.03 0.16
S4 DexMV [23] DAPG [19] 101 / 91 0.27 0.31
S7 Ours PPO [66] 1 / 1 0.97 0.68
V5 Ours Behavior cloning 1 / 1 0.97 0.68
IV-C3 Learning a unified multi-object visual policy from multiple videos
Finally, we train a single policy for all objects and configurations and report its performance in Table VI. We test the policy under three initial configurations for each object as the above paragraph. The visual policy V9 has the same architecture as the policy V8 in Table V, which uses 3D point clouds represented in the world coordinate frame. By comparing V9 and V8, we can see that learning a unified model is more challenging than training policies for each object separately. Therefore, we improve the point cloud representations in the visual policy. In V10, we transform the 3D point clouds to the target coordinate system. As result, the model becomes more aware of its target and improves the performance significantly, i.e., 79% to 94% under ${\rm SR_{10}}$. To further incorporate fine-grained hand-object interaction features, we additionally transform 3D points into six hand joint coordinate systems in V11 (i.e., five finger tips and the hand wrist). Benefiting from extracting rich hand-object interaction features, our final model achieves an average success rate of 97% under ${\rm SR_{10}}$ for the five objects under the challenging test scenario with three different configurations.
IV-D Evaluation of our policy on Pour and Place inside Tasks
We further evaluate our approach on the pour and place inside tasks. Table VII shows the quantitative results. Rows S1 to S4 are state-based policies reported in [23] where they use 91 videos and 101 videos for pour and place inside tasks respectively. While we only adopt one video in training, our state-based policy S7 achieves 97% and 68% success rates on the two tasks respectively, significantly outperforming 27% and 31% success rates using their best models. Benefiting from high-quality trajectories generated by our state-based policies, our visual policies V5 also achieve good performance on these two tasks. Figure 6 illustrates qualitative examples from our visual policy for these two tasks.
V Conclusion
We introduce ViViDex, a new framework to learn vision-based dexterous manipulation from human videos. To this end, we first extract reference hand and object trajectories from human videos and train state-based policies via RL to generate diverse physically plausible trajectories as the training data for the vision-based policy. Then, the visual policy takes robot proprioceptive states and 3D point clouds as inputs and further transforms 3D point clouds into different coordinate systems to achieve better performance. We conduct extensive ablations experiments to validate the effectiveness of our ViViDex system. We show that our visual policies can outperform state-of-the-art accuracy by a great margin on three dexterous manipulation tasks. Future work aims at leveraging internet videos for acquiring more general dexterous manipulation skills and investigating advanced 3D pose estimators.
Acknowledgement: This work was granted access to the HPC resources of IDRIS under the allocation AD011013147 made by GENCI. It was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR19-P3IA-0001 (PRAIRIE 3IA Institute) and the ANR project VideoPredict (ANR-21-FAI1-0002-01).

### v2

@@
 Abstract
-In this work, we aim to learn a unified vision-based policy for a multi-fingered robot hand to manipulate different objects in diverse poses.
-Though prior work has demonstrated that human videos can benefit policy learning, performance improvement has been limited by physically implausible trajectories extracted from videos.
+In this work, we aim to learn a unified vision-based policy for multi-fingered robot hands to manipulate a variety of objects in diverse poses.
+Though prior work has shown benefits of using human videos for policy learning, performance gains have been limited by the noise in estimated trajectories.
 Moreover, reliance on privileged object information such as ground-truth object states further limits the applicability in realistic scenarios.
@@
 We then rollout successful episodes from state-based policies and train a unified visual policy without using any privileged information.
-A coordinate transformation method is proposed to significantly boost the performance.
-We evaluate our method on three dexterous manipulation tasks and demonstrate a large improvement over state-of-the-art algorithms.
-The project page is available at https://zerchen.github.io/projects/vividex.html.
+We propose coordinate transformation to further enhance the visual point cloud representation, and compare behavior cloning and diffusion policy for the visual policy training.
+Experiments both in simulation and on the real robot demonstrate that ViViDex outperforms state-of-the-art approaches on three dexterous manipulation tasks.
+Qualitative results are available at: zerchen.github.io/projects/vividex.html.
 Figure 1: The overall framework of our method for learning dexterous manipulation skills from human videos.
-It consists of three modules: the extraction of the reference trajectory from human videos, the trajectory-guided state-based policy learning using RL, and the vision-based policy learning using behavior cloning.
+It consists of three steps: extraction of reference trajectories from human videos, trajectory-guided state-based policy learning using RL, and vision-based policy learning using either the behavior cloning or the 3D diffusion policy.
 I Introduction
-Humans possess a remarkable ability to manipulate diverse objects using their hands based on visual perception.
-Nevertheless, compared with parallel grippers [1, 2, 3, 4, 5, 6], achieving comparable levels of dexterous manipulation with multi-fingered robot hands remains an ongoing challenge [7].
-It requires to convert visual signals to high-dimensional control commands.
-In recent years, the advancement of deep learning (DL) [8] and reinforcement learning (RL) [9, 10] has led to significant progress in learning-based algorithms for dexterous manipulation such as in-hand rotation [11, 12, 13], solving Rubik’s cube [14] and playing the piano like humans [15, 16].
-However, it proves challenging to train RL policies as RL training heavily relies on intricate reward engineering [10, 17] and extensive computational resources [14].
+People possess a remarkable ability to manipulate objects effortlessly with their hands, guided by visual perception.
+Despite significant progress, replicating the dexterity of human hands with multi-fingered robot hands remains challenging [1, 2, 3, 4, 5].
+This is largely due to the complexity of translating visual signals into the high-dimensional control commands required for dexterous manipulation.
+Recent developments in deep learning (DL) [6] and reinforcement learning (RL) [7, 8] have enabled significant progress in learning-based algorithms for dexterous manipulation such as in-hand rotation [9, 10, 11], solving Rubik’s cube [12] and playing the piano [13, 14].
+However, it proves challenging to train RL policies as RL training heavily relies on intricate reward engineering [8] and extensive computational resources [12].
 Furthermore, unnatural behaviors with high rewards can emerge from RL training [15].
-To address the limitations of RL training, some previous works [19, 20, 21, 22] have turned to imitation learning using robot demonstrations collected by teleoperation.
-While these approaches enhance training efficiency, they demand substantial human efforts to collect diverse robot demonstration data and are, hence, difficult to scale.
-As human videos capturing hands manipulating objects are more abundant and easy to acquire, an increasing research attention has been paid to leveraging human videos for learning dexterous manipulation skills [23, 24, 25, 26, 27].
-For example, DexMV [23] proposes to automatically extract robot and object poses from human videos, and then employs the data to accelerate RL via demonstration augmented policy gradient (DAPG) [19].
-However, since the extracted robot and object poses are noisy, the method needs hundreds of human videos to learn the manipulation of a single object and still requires reward engineering for different tasks.
-Additionally, existing works [23, 25, 28] take advantage of privileged information of objects in policy learning such as known object CAD models and ground-truth object poses, which are non-trivial to obtain from raw visual sensors in real-world scenarios.
-In this work, we propose a new framework called ViViDex to leverage human Videos for Vision-based Dexterous manipulation policy learning.
-It consists of three modules as illustrated in Figure 1: reference trajectory extraction, trajectory guided state-based policy learning, and unified vision-based policy learning.
-We first obtain reference trajectories for hands and objects from human videos following prior work [23].
-Though such trajectories are noisy and not directly suitable for robot control, they provide examples of natural hand-object interactions that we use as guidance.
-Our second module refines and recovers physics properties of the reference trajectory.
-We use RL for this purpose by training a state-based policy for each reference trajectory.
-We design a novel reward to keep pose similarity in the reference trajectory while solving the task.
-We also augment the trajectory during training to generalize to a broader range of object poses beyond the pose in the reference trajectory.
-Finally, we rollout episodes from optimized state-based policies and distill successful episodes to a unified vision-based policy without any privileged object information.
-A coordinate transformation method is proposed to improve the vision-based policy, which transforms the input 3D point clouds into hand joints and the target coordinate systems to capture fine-grained interaction features and increase awareness of the target position.
-We validate the effectiveness of our system on three dexterous manipulation tasks relocation, pouring and placing inside.
-Our ViViDex approach significantly outperforms the state-of-the-art method DexMV [23] while using significantly less human demonstrations.
+To address the limitations of RL, some previous works [16, 17, 18, 19, 20, 21] have turned to imitation learning using robot demonstrations collected by teleoperation.
+While these approaches enhance training efficiency, they demand substantial human efforts to collect diverse robot demonstrations and are, hence, difficult to scale.
+As human videos capturing hands manipulating objects [22, 23] are abundant and easy to acquire, there is a growing trend in robotics research to utilize human video demonstrations for learning dexterous manipulation skills [24, 25, 26, 27, 28].
+For example, DexMV [24] proposes to extract robot and object poses from human videos, and then employs the data to accelerate RL training via demonstration augmented policy gradient [16].
+However, since the extracted robot and object poses are noisy, the method needs hundreds of human videos to learn the manipulation of a single object and requires reward engineering for different tasks.
+Moreover, existing methods [24, 26, 29] often leverage privileged information of objects in policy learning such as ground-truth object CAD models and object poses, which are non-trivial to obtain from visual sensor data in real-world scenarios.
+In this work, we propose ViViDex, a new framework making use of human Videos for Vision-based Dexterous manipulation.
+ViViDex consists of three modules as illustrated in Figure 1.
+We first obtain human hand and object trajectories from video demonstrations [24].
+Though such trajectories are noisy and not directly usable for robot control, they provide examples of natural hand-object interactions.
+Our second module then refines the reference trajectories to be physically plausible.
+We train a state-based policy with RL for each reference trajectory.
+A novel trajectory-guided reward is proposed to keep pose similarity in the reference trajectory while solving the task.
+We also augment trajectories during training to generalize to a broader range of object poses beyond the pose in the reference trajectory.
+Finally, we rollout episodes from optimized state-based policies and distill successful episodes to a unified vision-based policy using no privileged object information.
+To this end, we propose to improve visual representation by transforming input 3D point clouds into hand-centered coordinate systems.
+We evaluate ViViDex on three dexterous manipulation tasks relocation, pour and placing inside in simulation and demonstrate significant improvements over the state-of-the-art DexMV [24] while using significantly less human demonstrations.
+Our visual policy also achieves good performance both in simulation and on a real robot for both seen and unseen objects.
 To summarize, our contributions are three-fold:
@@
 $\bullet$ We improve the RL rewards for the state-based policy by imitating trajectories extracted from human videos and propose a novel model architecture for the vision-based policy.
-$\bullet$ Experiments on three dexterous manipulation tasks demonstrate the effectiveness of the ViViDex approach.
+$\bullet$ Extensive simulation and real robot experiments demonstrate the effectiveness of our proposed ViViDex approach.
 II Related Work
 Dexterous manipulation.
-The multi-fingered robotic hands empower robots to perform delicate operations on objects like humans.
-Previous works can be generally grouped into two categories: trajectory optimization methods [29, 30, 31, 32, 33] and data-driven learning methods [19, 23, 24, 28, 34, 35, 36, 37].
-Trajectory optimization methods usually require well-defined dynamic models for the robot and the manipulated object, which are usually not easy to obtain in practice.
-Data-driven methods instead train neural network policies for manipulation, allowing to learn generalizable skills.
-Most of the works [19, 28, 35, 38, 39, 40] use reinforcement learning (RL) algorithms to train policies.
+Multi-fingered robotic hands enable robots to perform delicate manipulation operations on objects.
+Previous work addressed dexterous manipulation by trajectory optimization [30, 31, 32, 33, 34, 35, 36] and data-driven learning methods [16, 24, 25, 29, 37, 38, 39, 40, 41, 42, 43].
+Trajectory optimization methods usually require well-defined dynamic models for the robot and the manipulated object, which are sometimes hard to obtain in practice.
+Data-driven methods instead directly train neural policies given robot data.
+Most of the works [44, 45, 16, 29, 38, 46, 47, 48] use RL algorithms to train policies.
 Nagabandi et al.
@@
 To accelerate the convergence of RL, Rajeswaran et al.
-[19] introduce the demo augmented policy gradient algorithm (DAPG) and train RL with expert demonstrations.
+[16] introduce demonstration augmented policy gradients (DAPG) and train RL with expert demonstrations.
 Instead of manually collecting expert demonstrations, ILAD [29] derives demonstration data via an affordance model [49] and trajectory optimization [50].
@@
 [46] decompose the dexterous grasping into two sub-tasks: the grasp proposal generation and the goal-conditioned policy training and largely improve the generalization of the learned policy.
-However, RL is data-inefficient and requires well-designed rewards, which often results in unrealistic object manipulation actions.
-To address these problems, more recent works [23, 25, 26, 27] attempt to learn policies using human demonstrations and combine imitation learning with RL.
+However, RL is data-inefficient, requires well-designed rewards, and often results in unrealistic object manipulation.
+In this work, we leverage human videos to improve RL training of state-based policies and then train a unified visual policy that enables object manipulation given 3D point cloud inputs.
 Learning robotic manipulation from human videos.
@@
 Equipping robots with the capacity to acquire manipulation skills by simply watching human videos has been an attractive direction [56, 57, 58, 59, 60].
-Some works [34, 53, 54, 55] propose to use human videos to pre-train a visual representation for the policy learning.
-R3M [34] proposes to pre-train CNNs with egocentric human videos [56] and perform the imitation learning based on the learned visual representations.
-Xiao et al.
-[53] and Radosavovic et al.
-[54] propose to leverage in-the-wild human videos to pre-train the vision transformer and demonstrate the improvement on different manipulation tasks.
-The recent work H-InDex [55] employs 3D human pose estimation from human videos as a proxy task to pre-train the visual representation for dexterous manipulation tasks.
-Another line of works [57, 58, 59, 60] focus on learning the reward function from human videos.
+One line of works [37, 61, 62, 63, 64] aims to learn generic visual representations from large-scale human video data [65] and then makes use of such pre-trained representations to learn control policies.
+Another line of works [66, 67, 68] focuses on learning reward functions from human videos.
 Chen et al.
@@
 [67] propose a task-agnostic reward function by using unlabeled human videos.
-Chane-Sane et al.
-[59] formulate the reward learning as a retrieval task and propose a data-driven approach to map human demonstrations to unseen robotic manipulation tasks.
-Xiong et al.
-[60] propose to translate human demonstration videos into robot demonstration videos and use the keypoint-based representation to define the distance metric for the policy execution.
 Some other works [24, 25, 26, 27, 69] explore how to explicitly leverage human videos to facilitate learning dexterous manipulation skills.
@@
 [25, 69] propose to utilize object affordances priors and human grasping priors to improve the dexterous grasping policy.
-Shaw et al.
-[26] propose to first pre-train the policy with visual, action and physical priors learned from human videos and finetune the policy with teleoperated demonstrations to make it generalize to different robotic tasks.
-DexMV [23] and DexRepNet [25] extracts demonstrations data from human videos and uses DAPG algorithm [19] to launch the RL training with demonstrations.
-In this work, not limited to the simple dexterous grasping action [24, 25, 35, 38, 61], we aim to solve more complex tasks by utilizing human videos.
-To additionally address limitations of previous works [23, 25, 28] that require object states (e.g., object poses) during the policy execution, we propose learning vision-based policies that achieve high manipulation accuracy while only using proprioceptive robot information and depth observations in the form of point clouds.
+DexMV [24] and DexRepNet [26] extract demonstrations data from human videos and use DAPG algorithm [16] to launch the RL training with demonstrations.
+PGDM [28] proposes to initialize the robot hand configuration at the given pre-grasp derived from human motions and enable efficient RL training.
+Compared to prior work [24, 26], our method allows to solve more complex manipulation tasks with fewer videos by using a unified trajectory guided reward learned from videos and automatic trajectory augmentation.
 III Method
-TABLE I: The state definition. The left part summarizes robot states. The right part defines object states.
-Parameters Descriptions Parameters Descriptions
-$\bm{q}_{r}\in\mathbb{R}^{30}$ robot motor positions $\bm{q}_{o}\in\mathbb{R}^{6}$ object canonical positions
-$\bm{\dot{q}}_{r}\in\mathbb{R}^{30}$ robot motor velocities $\bm{\dot{q}}_{o}\in\mathbb{R}^{6}$ object canonical velocities
-$\bm{x}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot joint positions $\bm{x}_{o}\in\mathbb{R}^{3}$ object center position
-$\bm{\theta}_{\text{r}}\in\mathbb{R}^{16\times 4}$ robot joint rotations $\bm{\theta}_{o}\in\mathbb{R}^{4}$ object global rotation
-$\bm{\dot{x}}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot linear velocities $\bm{\dot{x}}_{\text{o}}\in\mathbb{R}^{3}$ object linear velocity
-$\bm{\omega}_{\text{r}}\in\mathbb{R}^{16\times 3}$ robot angular velocities $\bm{\omega}_{\text{o}}\in\mathbb{R}^{3}$ object angular velocity
-$\bm{t}_{\text{r}}\in\mathbb{R}^{7\times 3}$ robot target translations $\bm{t}_{\text{o}}\in\mathbb{R}^{3}$ object target translations
 This work aims to train vision-based dexterous manipulation policies.
 However, jointly learning the visual representation and the control policy is challenging.
-Therefore, we first train the state-based policy and then distill the generated data to a vision-based policy.
+Therefore, we first train the state-based policy and then distill the learned policy to a vision-based policy.
 To train the state-based policy, as the pure RL training is ineffective [24, 26, 29], we extract reference trajectories from human videos and propose trajectory-guided RL to improve the performance.
-In the following, we first present details about the reference trajectory generation in Section III-A.
+In the following, we first present the reference trajectory generation in Section III-A.
 Then, we introduce our state-based policy learning algorithm in Section III-B.
-Finally, we describe how we distill multiple state-based policies into the visual policy and its model architecture in Section III-C.
+Finally, we describe the visual policy and its training in Section III-C.
 III-A Reference Trajectory Extraction
-In this work, we use the DexYCB [43] video dataset, which captures human hand grasping of objects.
-We first extract hand and object poses from the video.
-The hand poses and shapes are represented by MANO [62] and result in 3D locations of joints $\bm{\psi}_{h}\in\mathbb{R}^{21\times 3}$.
-Then, we aim to retarget the human hand pose to the robot hand pose.
-We use the Adroit anthropomorphic robot [19], which consists of 30 degrees of freedom $\bm{q}_{r}\in\mathbb{R}^{30}$.
+To guide our policy training, we extract hand and object poses from video demonstrations.
+The poses and shapes of hands are represented by MANO [70] and result in 3D locations of joints $\bm{\psi}_{h}\in\mathbb{R}^{21\times 3}$.
+To account for differences in geometry, we retarget the human hand pose to the robot hand pose.
 Following [24, 71, 72], we formulate the hand motion retargeting from a video of length $T$ as an optimization problem and define its objective function as:
-$\begin{split}&\min_{\bm{q}_{r}^{t}}\sum_{t=1}^{T}\Big{\|}\bm{\hat{x}}_{rt}^{t}(\bm{q}_{r}^{t})-\bm{{\psi}}_{ht}^{t}\Big{\|}^{2}_{2}\\ &+\Big{\|}\bm{\hat{x}}_{rp}^{t}(\bm{q}_{r}^{t})-\bm{{\psi}}_{hp}^{t}\Big{\|}^{2}_{2}+\alpha\Big{\|}\bm{q}_{r}^{t}-\bm{q}_{r}^{t-1}\Big{\|}^{2}_{2},\end{split}$ (1)
-where $\bm{{\psi}}_{ht}^{t}\in\mathbb{R}^{5\times 3}$ and $\bm{{\psi}}_{hp}^{t}\in\mathbb{R}^{5\times 3}$ denote human hand tip and middle phalanx positions, respectively.
-We solve their counterparts $\bm{\hat{x}}_{rt}^{t}$ and $\bm{\hat{x}}_{rp}^{t}$ in the robot via forward kinematics according to given ${\bm{q}_{r}^{t}}$.
-In the first two terms, we aim to minimize the difference between the 3D locations of robot joints to those in the human hand.
-We also use a regularization term to avoid sudden changes in the robot pose, where we empirically set $\alpha$ and ${\bm{q}_{r}^{0}}$ to 4e-3 and the mean pose of its motion limits, respectively.
-We use the NLopt solver [65] for the optimization and port robot and object motions into the simulator to generate reference trajectories, which are only visually plausible but not physically plausible.
-Figure 2: The illustration of different stages for the reference trajectory.
-The pre-grasp stage (a) describes how the robot approaches the object.
-The manipulation stage (b) focuses on how the robot hand manipulates the object.
+$\begin{split}\min_{\bm{q}_{r}^{t}}\sum_{t=1}^{T}\Big{\|}\bm{\hat{x}}_{rj}^{t}(\bm{q}_{r}^{t})-\bm{{\psi}}_{hj}^{t}\Big{\|}^{2}_{2}+\alpha\Big{\|}\bm{q}_{r}^{t}-\bm{q}_{r}^{t-1}\Big{\|}^{2}_{2},\end{split}$ (1)
+where $\bm{q}_{r}^{t}$ represents robot joint rotation angles.
+$\bm{{\psi}}_{hj}^{t}$ denotes human hand tip and middle phalanx positions.
+We solve their robot counterparts $\bm{\hat{x}}_{rj}^{t}$ via forward kinematics according to given ${\bm{q}_{r}^{t}}$.
+The first term aims to minimize the difference between the 3D locations of robot joints to those in the human hand.
+We also use a regularization term to avoid sudden changes in the robot pose, where we empirically set $\alpha$ to 4e-3 and ${\bm{q}_{r}^{0}}$ to the mean pose of its motion limits.
+We use the NLopt solver [73] for the optimization and port robot and object motions into the simulator to generate reference trajectories, which are visually plausible but not physically plausible.
+Figure 2 illustrates results of the motion retargeting for the Allegro robot hand.
+Figure 2: Motion retargeting results for the Allegro hand and objects under different poses for selected DexYCB videos.
 III-B State-based Policy
-We employ the generated reference trajectory as the reward function for RL to constrain the robot hand and object motions and train a state-based policy to recover physically plausible trajectories.
+We train a state-based policy with RL to recover physically plausible trajectories, where the above generated reference trajectories are employed in the reward function to guide robot hand and object motions.
 The network architecture for the state-based policy consists of actor and critic MLPs [74].
-It takes both robot and object states described in Table I as inputs and predicts the robot control commands.
-Different from DexMV [23] that needs different reward functions for different manipulation tasks, our method can solve different tasks with the same reward functions.
+It takes both robot and object states as inputs and predicts the robot control commands.
 In the following, we will describe our reward functions and the trajectory augmentation for RL.
-Reward functions.
-To define the reward function, we divide the reference trajectory into two stages shown in Figure 2: the pre-grasp stage and the manipulation stage.
-The recent work PGDM [27] focuses on the manipulation stage and initializes the robot hand at a given pre-grasp configuration for RL, which is a simplified setting compared with ours.
-During the pre-grasp stage, the robot hand approaches the object without making physical contacts.
-During RL training, we require that the robot approaches the object in a similar way as humans and define the following reward function:
+Trajectory-guided reward functions. During RL training, we propose to divide the reference trajectory into two stages: the pre-grasp stage and the manipulation stage.
+During the pre-grasp stage, the robot hand needs to approach the object without making physical contacts.
+We require that the robot approaches the object in a similar way as humans and define the following reward function as:
 $\begin{split}R_{p}=\sum_{t=1}^{T_{p}}10\cdot{\rm exp}(-10\cdot\Big{\|}\bm{{x}}_{rt}^{t}(\bm{q}_{r}^{t})-\bm{\hat{x}}_{rt}^{t}\Big{\|}^{2}_{2}),\end{split}$ (2)
-where $T_{p}$ is the length of pre-grasp steps, and $\bm{\hat{x}}_{rt}^{t}\in\mathbb{R}^{5\times 3}$ denotes the robot finger tip positions at the timestep $t$ in the reference trajectory $\bm{{x}}_{rt}^{t}$ is the current robot tip locations.
+where $T_{p}$ is the length of pre-grasp steps, and $\bm{\hat{x}}_{rt}^{t}$ denotes the robot finger tip positions at the timestep $t$ in the reference trajectory.
+$\bm{{x}}_{rt}^{t}$ is the current robot finger tip positions.
 When the robot successfully reaches its pre-grasp configuration, the episode starts its manipulation stage.
-In this stage, the robot manipulates the object and aims to bring it to the desired target configuration.
+In this stage, the robot aims to manipulate the object and bring it to the desired target configuration.
 Here, we define the reward function to constrain the robot and object motions jointly:
-$\begin{split}&R_{m}=\sum_{t=T_{p}+1}^{T_{r}}\lambda_{1}\cdot{\rm exp}(-\alpha_{1}\cdot\Big{\|}\bm{{x}}_{rt}^{t}(\bm{q}_{r}^{t})-\bm{\hat{x}}_{rt}^{t}\Big{\|}^{2}_{2})\\ &+\lambda_{2}\cdot{\rm exp}(-\alpha_{2}\cdot(\Big{\|}\bm{x}_{o}^{t}-\bm{\hat{x}}_{o}^{t}\Big{\|}^{2}_{2}+\alpha_{3}\cdot\phi(\bm{\theta}_{o}^{t},~{}\bm{\hat{\theta}}_{o}^{t})))\\ &+\sum_{i=1}^{5}\lambda_{3}\cdot\mathds{1}^{\rm contact}_{\rm finger~{}i}+\lambda_{4}\cdot\mathds{1}^{\rm lift}_{\rm object},\end{split}$ (3)
-where $T_{r}$ refers to the length of the whole reference trajectory.
-The first term of (3) constrains the hand motions similar to (2).
-The second term constrains the object motions, and $\bm{x}_{o}^{t}$ and $\bm{\hat{x}}_{o}^{t}$ represent the current object position and the reference object position at timestep $t$.
+$\begin{split}R_{m}=\sum_{t=T_{p}+1}^{T_{r}}\lambda_{1}R_{m}^{h}+\lambda_{2}R_{m}^{o}+\lambda_{3}\mathds{1}_{\rm cont}+\lambda_{4}\mathds{1}_{\rm lift},\end{split}$ (3)
+where $T_{r}$ is the length of the reference trajectory.
+The first term $R_{m}^{h}$ constrains the hand motions similar to (2).
+The second term $R_{m}^{o}={\rm exp}(-\alpha_{1}(\Big{\|}\bm{x}_{o}^{t}-\bm{\hat{x}}_{o}^{t}\Big{\|}^{2}_{2}+\alpha_{2}\phi(\bm{\theta}_{o}^{t},~{}\bm{\hat{\theta}}_{o}^{t})))$ constrains the object motions, where $\bm{x}_{o}^{t}$ and $\bm{\hat{x}}_{o}^{t}$ are the current object position and its reference at timestep $t$.
 $\phi(\cdot)$ computes the angular distance between the current object orientation with its reference $\bm{\hat{\theta}}_{o}^{t}$.
-In the third term, we encourage contacts between each finger of the robot and the object.
+The third term computes the number of finger tips in contact with the object.
 The forth term assigns bonus points when the object is lifted off the table.
-The second and forth terms are introduced in [27], and we use the same hyper-parameters (i.e., $\lambda_{2}=10$, $\alpha_{2}=50$, $\alpha_{3}=0.1$ and $\lambda_{4}=2.5$).
-For our proposed first and third terms, we empirically set $\lambda_{1}$, $\alpha_{1}$ and $\lambda_{3}$ to 4, 10 and 0.5, respectively.
-When the robot and object reaches their targets, we want them to remain stable in their target configurations for a while, which is especially important for the pour task to make sure that all particles in the mug have enough time to drop into the target container.
-To achieve this when $t~{}>~{}T_{r}$, we still use (3) and constrain the robot and object to stay in their final target configurations until the end of the episode.
-Figure 3: Illustration of our vision-based policy architecture.
+We empirically set $\lambda_{1}=4,\lambda_{2}=10,\lambda_{3}=0.5,\alpha_{1}=50$ and $\alpha_{2}=0.1$.
+Since our reward function is primarily derived from trajectories in videos, it can be applied across various manipulation tasks, reducing the need for task-specific reward engineering as in DexMV [24].
 Reference trajectory augmentation.
-By using our proposed reward functions, though the state-based policy can successfully imitate the hand and object motions from a video, it still cannot generalize to different initial object positions, rotations and target positions.
-To make our policy applicable to different initial object configurations and target positions, we introduce our reference trajectory augmentation strategies during RL training.
-To deal with different initial object positions or rotations, we could accordingly transform the whole reference trajectory.
-When we randomly set the target object position, we augment the original object trajectory through interpolations aimed at reaching the specified target position.
-For pour and place inside tasks, we first design the desired final object configuration.
-In the pour task, we ensure that the mug is tilted appropriately, which allows the particles inside to effortlessly drop into the target container.
-Our goal for place inside is to accurately position the banana within the container.
-Then, we interpolate the object position and orientation to reach the desired configuration.
-Similarly, we can also interpolate hand motions and use our proposed reward functions to optimize the policy.
+Though the state-based policy can successfully imitate the hand and object motions from a video, it remains challenging to generalize to different initial object positions, rotations and target positions.
+To make our policy applicable to different initial and target object configurations, we introduce our reference trajectory augmentation strategies during RL training.
+Specifically, we randomly set initial object positions or rotations and then transform the whole reference trajectory accordingly.
+In order to further increase the diversity of target object positions, we augment the original object trajectory through interpolations between the original last object pose and the target pose.
+We similarly interpolate hand motions and use proposed reward functions to train the state-based policy.
 III-C The Vision-based Policy
@@
 To generate training data for the visual policy, we rollout the optimized state-based policy and generate diverse physically plausible trajectories.
-During the rollout process, we compute the 3D point clouds ${{\rm PC_{w}}\in\mathbb{R}^{N\times 3}}$ from three depth cameras in the simulator, where $\rm PC_{w}$ is represented under the world coordinate system (i.e., the center of the table) and $N$ is the number of points.
-Figure 4: Motion retargeting results for selected DexYCB [43] videos. The three rows show different initial object poses for each object category.
-As shown in Figure 3, inspired by [5], we propose to transform $\rm PC_{w}$ into the desired target coordinate system $\rm PC_{t}$, which make the model more aware of the target position for control predictions.
-To capture dense interaction features between the robot and the object, we additionally transform $\rm PC_{w}$ into six different robot joint coordinate systems (i.e., one wrist and five fingers).
-Finally, we combine point clouds representations under different coordinate systems ${\rm PC}\in\mathbb{R}^{N\times 24}$ and feed them into PointNet [67].
-Our visual policy predicts control commands based on extracted visual features and robot states.
-We train our visual policy using behavior cloning from generated physically plausible trajectories.
+During the rollout process, we compute the 3D point clouds ${{\rm PC_{w}}\in\mathbb{R}^{N\times 3}}$ from the depth camera, where $\rm PC_{w}$ is represented under the world coordinate system (i.e., the center of the table) and $N$ is the number of points.
+Coordinate transformation.
+As shown in Figure 1, inspired by [75], we propose to transform $\rm PC_{w}$ into the desired target coordinate system $\rm PC_{t}$, which makes the model more aware of the target position for control predictions.
+To capture dense interaction features between the robot and the object, we additionally transform $\rm PC_{w}$ into different robot joint coordinate systems (i.e., palm and finger tips).
+Finally, we combine point clouds representations under different coordinate systems ${\rm PC}\in\mathbb{R}^{N\times 3(j+3)}$ and feed them into PointNet [76], where $j$ is the number of finger tips.
+Our visual policy predicts control commands based on extracted visual features and robot proprioceptive states.
+Training.
+We train our visual policy using either the behavior cloning (BC) or the recently proposed 3D diffusion policy [77, 78] from generated physically plausible trajectories.
+The BC model directly takes transformed point clouds and robot states as inputs and predicts robot actions.
+The diffusion policy employs the extracted 3D features from PointNet [76] as the global condition for the denoising model and recovers actions from Gaussian noise.
+We train both models with $\ell 2$ loss between the predicted and ground-truth actions.
 IV Experiments
-We conduct extensive experiments to evaluate our state-based and visual policies on three manipulation tasks, namely relocate, pour and place inside tasks.
+We conduct extensive experiments and evaluate state-based and visual policies learned by our approach on three manipulation tasks for five seen and ten unseen objects.
 IV-A Experimental Setting
 Video dataset.
-The DexYCB [43] dataset contains human demonstration videos of hand-object interactions for 20 YCB objects with 20 videos per object.
+The DexYCB [51] dataset contains human demonstration videos of hand-object interactions.
 Following [24], we focus on five objects including mustard bottle, tomato soup can, sugar box, large clamp and mug.
-Due to the unavailability of DexMV videos [23], we select videos from the DexYCB dataset that shares high visual similarities with DexMV videos.
-For each object, we choose one video which is most similar to those used in DexMV and use this video for most of our experiments; the top row in Figure 4 presents motion retargeting results for these videos.
-For the final experiment, we add additional initial configurations of the objects to the test set (different from [23]) and use two additional videos per object with different initial object positions for training, as illustrated in the last two rows of Figure 4.
-We use hand and object poses provided by the dataset to perform motion retargeting and generate reference trajectories.
-Simulation benchmark.
-We follow previous work [23, 27] and use the MuJoCo physics simulator [68] together with the Adroit robot hand [19].
-The robot hand consists of 30 degrees of freedom (DoF).
-We set the control frequency and simulation frequency as 40HZ and 400HZ, respectively.
-As discussed above, we use the same testing configuration as in [23] except for the final experiment, where we use a more difficult setting, and evaluate the performance of our method on three manipulation tasks:
-TABLE II: Performance of our state-based policy using hand rewards at different stages.
-It follows the human hand trajectory extracted from videos to manipulate objects (R1), reach the pre-grasp configurations (R2) or combine both (R3).
+For each object, we choose three videos which are most similar to those used in DexMV [^1] The videos in DexMV are not publicly released.
+in our experiments.
+Figure 2 presents motion retargeting results for these videos.
+Two evaluation protocols are used for benchmarking the performance of policies.
+Protocol #1 trains the policy for each object separately and uses initial object poses from the first row of Figure 2.
+For Protocol #2, we train a unified policy for all five objects and adopt three poses for each object shown in Figure 2.
+Furthermore, we evaluate the performance on ten unseen objects, namely master chef can, tuna fish can, pudding box, gelatin box, potted meat can, banana, pitcher base, bleach cleanser, wood block and foam brick.
+Simulation environment.
+We follow previous works [16, 24, 34] by using the the Adroit robot hand and the MuJoCo simulator for fair comparison.
+However, this setup is less realistic, as the robotic hand is not attached to an arm and can move freely.
+To address this, we attach an Allegro robot hand to a UR5 arm, mirroring our actual hardware configuration.
+We use the SAPIEN [79] simulator for this setup, which simplifies creating this more realistic environment and allows to speed up simulation.
+Unless specified otherwise, the Adroit hand in MuJoCo is used for benchmarking against state-of-the-art methods, while the Allegro hand in SAPIEN is employed for training visual policies and performing ablation studies.
+Manipulation tasks.
+We follow [24] to evaluate on three tasks.
+The first relocate task requires the robot to move an object to a target position.
+The pour task aims to grasp a mug filled with particles and pour the particles into a container.
+The success rate is measured by the percentage of particles in the container.
+The place inside task aims to grasp a banana and place it into a mug.
+The success rate is measured by the percentage of the banana mesh in the mug.
+Figure 3: Qualitative comparison of state-based policies using different rewards for Protocol #1 and the Allegro hand.
+R1 (w/o hand reward in pre-grasp) leads to unstable grasps.
+R2 (w/o hand reward in manipulation) results in unnatural hand actions.
+Our proposed approach R3 uses hand rewards at both stages and achieves the best performance.
+TABLE I: Performance of our state-based policy using different hand rewards (Allegro hand, Protocol #1).
+It follows the extracted human hand trajectory to manipulate objects (R1), reach the pre-grasp location (R2) or combine both (R3).
 Hand Reward ${\rm E_{o}}\downarrow$ ${\rm E_{h}}\downarrow$ ${\rm SR_{o}}\uparrow$ ${\rm SR_{h}}\uparrow$ ${\rm SR_{3}\uparrow}$
 Pre-grasp Manipulation
-R1 $\times$ 0.063 0.15 0.37 0.00 0.00
-R2 $\times$ 0.0031 0.053 0.93 0.32 1.00
-R3 0.0031 0.024 0.95 0.83 1.00
-Relocate task requires the robot to move an object to a target position. The task is successfully solved when the object reaches the target position within a certain threshold.
-Pour task requires the robot to grasp a mug filled with particles and pour the particles into a container. The success rate is measured by the percentage of particles in the container.
-Place inside task requires the robot to grasp a banana and place it into a mug. The success rate is measured by the percentage of the banana mesh in the mug.
+R1 $\times$ 0.048 0.21 0.35 0.00 0.00
+R2 $\times$ 0.0033 0.077 0.95 0.32 1.00
+R3 0.0019 0.032 0.97 0.79 1.00
 Evaluation metrics.
@@
 To measure the accuracy more precisely, we add a more rigorous threshold of 3cm and compute the success rate ${\rm SR_{3}}$.
-For the state-based policy, we additionally evaluate how well the learned physically plausible trajectory matches the reference trajectory in the video using four metrics [27]: ${\rm E_{o}}$ computes the average object position error between the current object trajectory and its reference trajectory over time; ${\rm E_{h}}$ computes the average finger tips position error between the current robot hand trajectory and its reference trajectory over time; ${\rm SR_{o}}$ reports the fraction of timesteps where ${\rm E_{o}}$ is below the 1cm threshold; ${\rm SR_{h}}$ computes the fraction of timesteps where ${\rm E_{h}}$ is lower than 5cm.
+For the state-based policy, we additionally evaluate how well the physically plausible trajectory matches the reference trajectory using four metrics [28]: ${\rm E_{o}}$ computes the average object position error between the current object trajectory and its reference trajectory over time; ${\rm E_{h}}$ computes the average finger tips position error between the current robot hand trajectory and its counterpart over time; ${\rm SR_{o}}$ reports the fraction of timesteps where ${\rm E_{o}}$ is below 1cm; ${\rm SR_{h}}$ computes the fraction of timesteps where ${\rm E_{h}}$ is lower than 5cm.
 Implementation details.
-During RL training, we randomly set the object initial and target positions following [23] and also randomize the initial object orientation.
-During testing, we run 300 episodes with different initial configurations and report the average success rate.
-For the state-based policy, we use both robot and object states defined in Table I as inputs and predict the control command.
+We use PPO [74] to optimize the state-based policy.
+The RL training takes around 2 hours on a single A100 GPU for each video.
+We rollout 100 successful trajectories from the state-based policy for each video, where trajectories differ in initial object positions, orientations and target positions.
+During rollouts, we render 3D scene point clouds from the depth camera.
+These data are used to train our visual policy, which takes robot joint positions and 3D point clouds as inputs.
+Training our visual policy on a single A100 GPU takes around 10 hours using behavior cloning and about 20 hours using the diffusion policy, both based on data from 15 videos.
+For testing, we run 300 episodes with different initial configurations and report the average success rate.
 The length of a single episode is 60, 80 and 100 for the relocate task, place inside task and pour task, respectively.
-We use PPO [66] to optimize the state-based policy.
-We set the learning rate and batch size to 1e-5 and 256, respectively.
-The RL training takes around 30 hours on a single V100 GPU and 32 CPUs for each video separately.
-Our visual policy only requires robot proprioceptive states (the left part of Table I) and point clouds as inputs.
-To train the visual policy, we rollout 500 successful trajectories from the state-based policy for each video, where trajectories differ in initial object positions, rotations and target positions.
-During rollouts, we render 3D scene point clouds from three depth cameras surrounding the table in the simulation platform.
-We train the visual policy with Adam optimizer using a learning rate of 1e-5 and a batch size of 64.It takes 8 hours on 4 A100 GPUs for our visual policy trained from 5 videos.
+The real robot takes around 1 minute to execute an episode.
 IV-B Evaluation of the State-based Policy on the Relocate Task
-We adopt the relocate task to evaluate the state-based policy.
-Specifically, we use one DexYCB video for each of the five object categories and train a state-based policy for each video separately.
-The initial object poses of the videos are illustrated in the first row of Figure 4.
-TABLE III: Comparison with state-of-the-art methods on the relocate task for state-based policies.
-We optimize a state-based policy using a single human video for each object (S6, S7), while previous methods usually need about one hundred human videos for each object (S2-S4).
+We evaluate the state-based policy on the relocate task using both the Adroit and Allegro hands.
+We follow Protocol #1 to train the state-based policy then evaluate each policy on the same object with novel placement of the object.
+Specifically, we randomly change the initial object position and rotation around z-axis for each testing episodes.
+TABLE II: Comparison with state-of-the-art methods on relocate task for state-based policies.
+We optimize a state-based policy using a single human video for each object (S6, S7, S8) (Protocol #1), while previous methods usually need about 100 videos for each object (S2-S4).
 * indicates that we re-train the method from [24] with 20 DexYCB videos for each object.
-Methods Training algorithms Num. videos Mustard bottle Tomato can Sugar box Large clamp Mug Avg.
+Methods Robot Hands Training algorithms Num. videos Mustard bottle Tomato can Sugar box Large clamp Mug Avg.
 ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
-S1 DexMV [23] TRPO [69] 0 0.06 - 0.67 - 0.00 - 0.51 - 0.49 - 0.35 -
-S2 DexMV [23] SOIL [70] 97 0.33 - 0.98 - 0.67 - 0.89 - 0.71 - 0.72 -
-S3 DexMV [23] GAIL+ [71] 97 0.06 - 0.66 - 0.00 - 0.52 - 0.53 - 0.50 -
-S4 DexMV [23] DAPG [19] 97 0.93 - 1.00 - 0.00 - 1.00 - 1.00 - 0.79 -
-S5 DexMV [23] DAPG [19] 20 1.00 0.63 1.00 0.56 0.09 0.00 0.08 0.00 0.02 0.00 0.44 0.24
-S6 Ours / wo rot. PPO [66] 1 0.75 0.03 0.74 0.65 0.38 0.08 0.39 0.03 0.98 0.03 0.65 0.16
-S7 Ours PPO [66] 1 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
-TABLE IV: Performance of separate visual policies for each object on the relocate task.
-The initial configurations in testing are derived from the first row of Figure 4 for each object.
-Our visual policies (V4-V6) only utilize one video in training.
-“Comp.
-PC” denotes complete point clouds sampled from object CAD models without occlusion.
-Source Models Comp. PC Object states Visual inputs Num. Points Noisy Points Mustard bottle Tomato can Sugar box Large clamp Mug Avg.
-${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
-V1 S4 512 $\times$ 1.00 0.67 1.00 0.50 0.14 0.00 1.00 0.77 1.00 0.58 0.83 0.50
-V2 S4 $\times$ 512 $\times$ 1.00 0.62 1.00 0.48 0.14 0.00 1.00 0.76 0.99 0.55 0.83 0.48
-V3 S4 $\times$ $\times$ 512 $\times$ 0.98 0.55 1.00 0.46 0.14 0.00 1.00 0.76 0.99 0.54 0.82 0.46
-V4 S7 $\times$ $\times$ 512 $\times$ 1.00 0.89 1.00 0.86 1.00 0.96 1.00 0.91 1.00 1.00 1.00 0.92
-V5 S7 $\times$ $\times$ 2048 $\times$ 1.00 0.97 1.00 0.80 1.00 0.98 1.00 0.93 1.00 1.00 1.00 0.94
-V6 S7 $\times$ $\times$ 2048 0.97 0.77 0.97 0.57 1.00 0.88 1.00 0.89 1.00 1.00 0.99 0.82
-Figure 5: Qualitative comparison of state-based policies using different rewards.
-R1 (w/o hand reward in pre-grasp) leads to unstable grasps.
-R2 (w/o hand reward in manipulation) results in unnatural hand actions.
-Our proposed policy R3 uses hand rewards at both stages and achieves the best performance.
+S1 DexMV [24] Adroit TRPO [80] 0 0.06 - 0.67 - 0.00 - 0.51 - 0.49 - 0.35 -
+S2 DexMV [24] Adroit SOIL [81] 97 0.33 - 0.98 - 0.67 - 0.89 - 0.71 - 0.72 -
+S3 DexMV [24] Adroit GAIL+ [82] 97 0.06 - 0.66 - 0.00 - 0.52 - 0.53 - 0.50 -
+S4 DexMV [24] Adroit DAPG [16] 97 0.93 - 1.00 - 0.00 - 1.00 - 1.00 - 0.79 -
+S5 DexMV [24] Adroit DAPG [16] 20 1.00 0.63 1.00 0.56 0.09 0.00 0.08 0.00 0.02 0.00 0.44 0.24
+S6 Ours Adroit PPO [74] 1 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
+S7 Ours Allegro PPO [74] 1 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
+S8 Ours w/o rot. Allegro PPO [74] 1 0.85 0.85 1.00 0.99 0.99 0.97 0.98 0.95 0.97 0.94 0.96 0.94
+TABLE III: Performance of visual policies for each object on relocate task with ${\rm SR_{3}}$ using Protocol #1 and Allegro hand.
+Algorithms #Points Mb. Tc. Sb. Lc. Mug Avg.
+V1 BC 512 0.99 0.82 0.82 0.78 0.90 0.86
+V2 BC 2048 1.00 1.00 0.87 0.91 1.00 0.96
+V3 Diffusion 2048 1.00 0.97 0.98 0.99 0.99 0.99
 IV-B1 Ablation Studies
 We first perform ablations to validate the importance of our hand rewards at the pre-grasp and the manipulation stages in state-based policy learning.
-Table II presents the averaged performance over the five videos for policies with different reward functions.
-The policy in R1 uses the proposed hand rewards (i.e., Equation 3) in the manipulation stage but follows the previous reward function [23, 28, 19] in the pre-grasp stage which is to minimize the distance between the robot hand and the object.
-It achieves poor performance due to the lack of guidance from human hand trajectory to approach the object.
-As shown in the first column of Figure 5, the policy is difficult to arrive at a plausible pre-grasp configuration to stably lift the object.
-The policy in R2 employs the proposed reward function in the pre-grasping stage (Equation 2) but follows PGDM [27] (i.e., the second and forth terms in Equation 3) in the object manipulation stage.
-The reward function in PGDM constrains the object motions to be close to the reference trajectory but does not add any constraints on the hand motions, which results in unnatural robot hand actions as shown in the second column of Figure 5.
-Our proposed policy in R3 utilizes hand rewards in both stages, which outperforms R1 and R2 on all the metrics and produces robust and realistic grasps as shown in the last column of Figure 5.
+Table I presents the averaged performance over all objects for policies with different reward functions for the Allegro hand.
+The policy in R1 uses the proposed hand reward $R_{m}^{h}$ in the manipulation stage but follows the previous reward function [24, 29, 16] in the pre-grasp stage which is to minimize the distance between the robot and the object.
+It achieves poor performance due to the lack of guidance from the human hand trajectory to approach the object.
+As shown in the first column of Figure 3, the policy has difficulties in arriving at a plausible pre-grasp configuration to stably lift the object.
+The policy in R2 employs the proposed reward function $R_{p}$ in the pre-grasping stage but follows PGDM [28] (i.e., $R_{m}^{o}$ and $\mathds{1}_{\rm lift}$) in the manipulation stage.
+The reward function in PGDM constrains the object motions to be close to the reference trajectory but does not add any constraints on the hand motions, which results in unnatural robot actions as shown in the second column of Figure 3.
+Our proposed policy in R3 utilizes hand rewards in both stages, which outperforms R1 and R2 on all the metrics and produces more robust and realistic grasps as shown in the last column of Figure 3.
 IV-B2 State-of-the-art comparison
-Table III presents the quantitative comparison of our state-based policy and state-of-the-art models [23].
-S1 to S4 denote four model variants reported in [23].
-As their training data is not released, we re-train their best model DAPG [19] (S4) using DexYCB videos and report the results in S5 for more fair comparison.
+Table II presents the quantitative comparison of our state-based policy and state-of-the-art models [24].
+S1 to S4 denote four model variants reported in [24] for the Adroit hand.
+As their training data is not released, we re-train their best model DAPG [16] (S4) using 20 DexYCB videos and report the results in S5 for fair comparison.
 We can see that the performance drops significantly for some objects, i.e., large clamp and mug.
-This indicates that their method requires sufficient amount of videos to learn the manipulation of objects with complex shapes.
-Our proposed policy in S7 significantly outperforms the previous methods despite only using one training video per object.
+This suggests that their approach needs a large number of videos to effectively learn how to handle objects with complex shapes.
+Our proposed policy in S6 significantly outperforms all previous methods despite only using one training video per object.
 Though S4 also performs well for mustard bottle, tomato soup can, large clamp and mug, it easily knocks down thin objects (e.g., sugar box) when the robot approaches the object and achieves poor performance.
 Our hand reward addresses this problem by mimicking the human pre-grasping trajectory.
-S6 is a variant of our model without rotation augmentation in training.
-The inferior performance of S6 to our final model S7 demonstrates the importance of augmentation to help improve the robustness of trained policies.
+As shown in S7, we achieve the same performance with the Allegro hand.
+S8 is a variant of our model S7 without rotation augmentation in the z-axis during training.
+The performance only drops insignificantly when tested for such rotations.
+This demonstrates the robustness of our method.
 IV-C Evaluation of the Visual Policy on the Relocate Task
@@
 IV-C1 Learning from a single video per object
-We first explore training separate visual policies for each object with only one video as in the previous section IV-B.
-During testing, we use the same object configurations as in the section IV-B, which are derived from object poses in the first row of Figure 4 by randomly changing the initial object position and rotation around z-axis for each testing episodes.
-The results are presented in Table IV.
-From V1 to V3, we rollout trajectories from the S4 models in Table III  [^1] We use S4 models available at released codes of DexMV [23].
-to train the visual policy.
-V1 has the same assumption as [28] which assumes known object states and complete object point clouds.
-V2 removes the dependency on complete point clouds but still requires object states.
-V3 further relaxes the assumption and only uses robot states together with scene point clouds.
-We can see that performance decreases when using no privileged object information (i.e., object states and complete object point clouds).
-From V4 to V6, we train the visual policy using rollouts from our state-based policies S7.
-Benefiting from high-quality trajectories generated by S7, our visual policy V4 can outperform V1 to V3 by a large margin without privileged information.
-To investigate how the density of point clouds affect the visual policy, we sample 2048 points in V5 at each timestep.
-This marginally improves the performance compared to V4.
-In V6 we evaluate the robustness of our visual policy by adding Gaussian noise with $\sigma$=3cm to 3D point clouds during training and testing.
-Our visual policy still functions reasonably under noisy point clouds observations.
-TABLE V: Performance of separate policies for each object on the relocate task.
-The initial configurations in testing are derived from all three rows in Figure 4.
-We compare our visual policies using different number of training videos.
-“Mb.”, “Tc.”, “Sb.” and “Lc.” denote mustard bottle, tomato can, sugar box and large clamp respectively.
-Num. videos Mb. Tc. Sb. Lc. Mug Avg.
-${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
-S4 97 0.46 0.28 0.59 0.35 0.11 0.00 0.63 0.31 0.65 0.62 0.49 0.31
-V5 1 0.33 0.32 0.41 0.25 0.33 0.32 0.43 0.31 0.58 0.57 0.42 0.35
-V7 2 0.65 0.42 0.66 0.60 0.66 0.62 0.67 0.64 0.91 0.85 0.71 0.63
-V8 3 0.96 0.66 0.95 0.82 0.89 0.63 0.98 0.92 0.98 0.97 0.95 0.80
-TABLE VI: Performance of unified visual policies for all objects on the relocate task.
-The initial configurations in testing are derived from all three rows in Figure 4.
-We compare point cloud representations with different coordinate systems where “wo.”, “ta.” and “ha.” represent world, target and robot hand.
-“Mb.”, “Tc.”, “Sb.” and “Lc.” denote mustard bottle, tomato can, sugar box and large clamp respectively.
-Coor. sys Mb. Tc. Sb. Lc. Mug Avg.
-wo. ta. ha. ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
-V9 $\times$ $\times$ 0.77 0.32 0.73 0.44 0.80 0.29 0.87 0.40 0.79 0.58 0.79 0.41
-V10 $\times$ 0.91 0.84 0.94 0.92 0.99 0.97 0.96 0.93 0.92 0.92 0.94 0.92
-V11 0.94 0.86 0.99 0.98 0.98 0.93 0.99 0.97 0.96 0.95 0.97 0.94
-IV-C2 Learning from multiple videos per object
-We next introduce a more challenging testing scenario where we evaluate policies under three different initial object configurations for each object as shown in Figure 4.
-For example, the mustard bottle does not necessarily stand on the table, but can lie in different ways.
-Table V presents the results of different models.
-The state-based policy S4 [23] in Table III though employs privilege object information and multiple videos per object in training, still generalizes poorly to different initial object poses.
-Our visual policy V5 in Table IV also suffers in this challenging testing scenario due to the insufficient training videos.
-From V7 to V8, we gradually increase the number of training videos with diverse initial object configurations and demonstrate significant increase in performance.
-Figure 6: Qualitative illustration of the performance of our visual policies on pour (a) and place inside (b) tasks.
-TABLE VII: Comparison with state-of-the-art methods on the pour and the place inside tasks for our policy models.
+We first investigate whether replacing the ground-truth object state in the state-based policy with visual point cloud inputs can affect the performance.
+For each video, we follow Protocol #1 to train a separate visual policy using the rollout data obtained from the state-based policy S7 in Table II, and evaluate under novel object placements as in the evaluation of the state-based policy.
+The results are presented in Table III.
+Compared with V1 that uses 512 points, V2 samples 2048 points and improves the overall performance, which shows the importance of point clouds densities.
+By comparing V2 and V3, the 3D diffusion policy [78] can achieve more robust performance than the plain behavior cloning algorithm.
+Our visual policies only perform slightly worse than our state-based policies.
+TABLE IV: Performance of unified visual policies on relocate task under $\rm{SR}_{3}$ using Protocol #2 and Allegro hand.
+We compare the performance using different number of human videos and point clouds under different coordinate systems.
+Algo. Vid. Coor. sys. Seen Unseen avg.
+target hand Mb. Tc. Sb. Lc. Mug Avg.
+V4 BC 1$\times$5 $\times$ $\times$ 0.33 0.33 0.33 0.30 0.42 0.34 0.31
+V5 BC 2$\times$5 $\times$ $\times$ 0.55 0.57 0.36 0.56 0.66 0.54 0.33
+V6 BC 3$\times$5 $\times$ $\times$ 0.69 0.80 0.83 0.86 0.88 0.81 0.38
+V7 BC 3$\times$5 $\times$ 0.93 0.92 0.99 0.93 0.99 0.95 0.37
+V8 BC 3$\times$5 0.93 0.97 0.97 0.98 1.00 0.97 0.41
+V9 Diff. 3$\times$5 0.97 0.99 1.00 0.99 0.99 0.99 0.50
+TABLE V: Comparison with state-of-the-art methods on the pour and the place inside tasks using the Adroit hand.
 Our approach uses a single human video for each task, while previous methods need more than 91 human videos.
 Methods Training algorithms Num. videos Pour success rate Place inside success rate
-S1 DexMV [23] TRPO [69] 0 / 0 0.01 0.03
-S2 DexMV [23] SOIL [70] 101 / 91 0.04 0.28
-S3 DexMV [23] GAIL+ [71] 101 / 91 0.03 0.16
-S4 DexMV [23] DAPG [19] 101 / 91 0.27 0.31
-S7 Ours PPO [66] 1 / 1 0.97 0.68
-V5 Ours Behavior cloning 1 / 1 0.97 0.68
-IV-C3 Learning a unified multi-object visual policy from multiple videos
-Finally, we train a single policy for all objects and configurations and report its performance in Table VI.
-We test the policy under three initial configurations for each object as the above paragraph.
-The visual policy V9 has the same architecture as the policy V8 in Table V, which uses 3D point clouds represented in the world coordinate frame.
-By comparing V9 and V8, we can see that learning a unified model is more challenging than training policies for each object separately.
-Therefore, we improve the point cloud representations in the visual policy.
-In V10, we transform the 3D point clouds to the target coordinate system.
-As result, the model becomes more aware of its target and improves the performance significantly, i.e., 79% to 94% under ${\rm SR_{10}}$.
-To further incorporate fine-grained hand-object interaction features, we additionally transform 3D points into six hand joint coordinate systems in V11 (i.e., five finger tips and the hand wrist).
-Benefiting from extracting rich hand-object interaction features, our final model achieves an average success rate of 97% under ${\rm SR_{10}}$ for the five objects under the challenging test scenario with three different configurations.
+L1 DexMV [24] TRPO [80] 0 / 0 0.01 0.03
+L2 DexMV [24] SOIL [81] 101 / 91 0.04 0.28
+L3 DexMV [24] GAIL+ [82] 101 / 91 0.03 0.16
+L4 DexMV [24] DAPG [16] 101 / 91 0.27 0.31
+L5 Ours PPO [74] 1 / 1 0.97 0.68
+L6 Ours Diffusion 1 / 1 0.97 0.68
+IV-C2 Learning a unified multi-object visual policy from multiple videos
+Finally, we adopt Protocol #2 and train a single policy for all objects with the Allegro hand and report its performance in Table IV.
+We test the policy under three initial poses for each object shown in Figure 2.
+From V4 to V6, we gradually increase the number of videos for each object from one to three and observe a significant improvement in performance.
+Then, V7 additionally transforms the 3D point clouds into the target coordinate system.
+As a result, the model becomes more aware of its target and largely improves the average performance for seen objects from 81% to 95%.
+To further incorporate fine-grained hand-object interaction features, we further transform 3D points into hand joint coordinate systems in V8.
+Benefiting from extracting rich hand-object interaction features, V8 achieves even better performance than V7.
+Different from V8, V9 trains the visual model using 3D diffusion policy and achieves an average success rate of 99% under the challenging test scenario.
+To investigate the generalization of our visual policies, we evaluate their performance using three different initial poses for ten unseen YCB objects.
+In Figure IV from V4 to V6, our model better generalizes to novel objects when using more videos.
+V8 incorporates fine-grained interaction features and enhances generalization.
+Compared with the behavior cloning, V9 significantly improves the unseen success rate from 41% to 50% by using the 3D diffusion policy, which demonstrates better robustness and generalization abilities.
 IV-D Evaluation of our policy on Pour and Place inside Tasks
-We further evaluate our approach on the pour and place inside tasks.
-Table VII shows the quantitative results.
-Rows S1 to S4 are state-based policies reported in [23] where they use 91 videos and 101 videos for pour and place inside tasks respectively.
-While we only adopt one video in training, our state-based policy S7 achieves 97% and 68% success rates on the two tasks respectively, significantly outperforming 27% and 31% success rates using their best models.
-Benefiting from high-quality trajectories generated by our state-based policies, our visual policies V5 also achieve good performance on these two tasks.
-Figure 6 illustrates qualitative examples from our visual policy for these two tasks.
+We further evaluate our approach on the pour and place inside tasks and report results in Table V.
+Rows L1-L4 show results for state-based policies [24] using the Adroit hand as well as 91 and 101 video demonstrations for pour and place inside tasks respectively.
+While we use only single human video for training, our state-based policy L5 for the Adroit hand achieves 97% and 68% success rates on these two tasks respectively, significantly outperforming 27% and 31% success rates corresponding to the best models of [24].
+By taking advantage of high-quality trajectories generated by L5, our visual policy L6 also achieves high performance.
+TABLE VI: Success rate of visual policies on relocate task using the real robot. We evaluate 10 episodes for each object.
+Methods Unified policy Seen Unseen Avg.
+Mb. Tc. Sb. Lc. Mug Avg.
+R1 BC $\times$ 10/10 8/10 9/10 8/10 9/10 0.88 -
+R2 BC 9/10 7/10 7/10 5/10 8/10 0.72 0.58
+R3 Diffusion 10/10 7/10 8/10 7/10 8/10 0.80 0.68
+Figure 4: Illustrations of our real-world robot experimental setup and the performance of our proposed ViViDex algorithm.
+IV-E Real world experiments
+To further demonstrate the advantages of ViViDex, we evaluate its performance on real-world dexterous manipulation.
+As shown in Figure 4(a), we use a UR5 robotic arm equipped with an Allegro hand and a single RealSense D435 RGB-D camera.
+Since point clouds for the real-world experiments are noisier than those in simulation, we use the state-based policy to collect data in the real world and then learn a unified visual policy based on real data.
+We run state-based policies in the simulator for specific object locations and run them on the real robot by placing the object in the same locations.
+This allows us to collect 3D scene point clouds and robot states for each execution step and build the real-robot training dataset.
+We collect five robot trajectories for each object and train the visual policy using our collected real-robot data.
+As shown in Figure 4(b), our visual policy can then manipulate the object.
+We summarize the quantitative results on the real robot in Table VI.
+We adopt the initial pose for seen objects from the first row of Figure 2 and include five unseen objects for evaluation: cracker box, spray bottle, bleach cleanser, water bottle and pudding box.
+R1 trains five policies separately and achieve an average success rate of 88%.
+R2 and R3 learn unified policies for five objects and demonstrate the ability to grasp unseen objects.
+The diffusion policy R3 also achieves better performance than BC in real experiments.
 V Conclusion
-We introduce ViViDex, a new framework to learn vision-based dexterous manipulation from human videos.
-To this end, we first extract reference hand and object trajectories from human videos and train state-based policies via RL to generate diverse physically plausible trajectories as the training data for the vision-based policy.
-Then, the visual policy takes robot proprioceptive states and 3D point clouds as inputs and further transforms 3D point clouds into different coordinate systems to achieve better performance.
-We conduct extensive ablations experiments to validate the effectiveness of our ViViDex system.
-We show that our visual policies can outperform state-of-the-art accuracy by a great margin on three dexterous manipulation tasks.
-Future work aims at leveraging internet videos for acquiring more general dexterous manipulation skills and investigating advanced 3D pose estimators.
-Acknowledgement: This work was granted access to the HPC resources of IDRIS under the allocation AD011013147 made by GENCI.
+We introduce ViViDex, a new framework for learning vision-based dexterous manipulation from human videos.
+Our approach extracts reference trajectories from human videos and uses them as a reward in training state-based policies with RL.
+This generates diverse physically plausible trajectories.
+We rollout these trajectories for training a visual policy, which takes as inputs robot proprioceptive states and 3D point clouds.
+To enhence visual policies, we further transforms 3D point clouds into hand-centered coordinate systems.
+We conduct extensive ablation experiments to validate the effectiveness of ViViDex on different simulators and a real robot.
+We show that our visual policies outperform state-of-the-art accuracy on different tasks by a significant margin and generalize to unseen objects.
+Future work aims at leveraging internet videos for acquiring more general dexterous manipulation skills and investigating advanced 3D pose estimation algorithms.
+Acknowledgement: We thank Ricardo Garcia Pinel for help with the camera calibration, Yuzhe Qin for clarifications about DexMV and suggestions on the real robot depolyment.
+This work was granted access to the HPC resources of IDRIS under the allocation AD011013147 made by GENCI.
 It was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR19-P3IA-0001 (PRAIRIE 3IA Institute) and the ANR project VideoPredict (ANR-21-FAI1-0002-01).